40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed May 5 18:14:42 2010, Dmass, Laser, Doubling, Massive Update plottedApr20Phase.pdf
    Reply  Wed May 19 19:33:55 2010, Dmass, Laser, Doubling, More Update Plotted14Phase.pdfplotted16RIN.pdfPlotted14Volts.pdfplotted16Volts.pdf
    Reply  Wed May 19 19:54:40 2010, Dmass, Laser, Doubling, Massive Update MZPhaseLowAvg.pdf
Message ID: 750     Entry time: Wed May 5 18:14:42 2010     Reply to this: 778   779
Author: Dmass 
Type: Laser 
Category: Doubling 
Subject: Massive Update 

There have been no updates since this post, I will remedy that. Hyperlinks to later more detailed elog posts will be added as they pop up.

Summary of situation at time of above post (When Matt Evans, Hartmut, Koji and I had a work party): This was around the time of the first big frontend crash, so we went totally analog

  • Locked PMC with an SR560 (30Hz pole and ~1000 gain)
  • Added locking to the Mach Zehnder via stuffable breadboard and PZT
  • Changed ISS to use the PD SUM as the error signal (because there seemed to be beam path induced sensing noise from using the PMC Trans)
  • Calibrated the Mach Zehnder via the Lissajous of the PD difference signals at each color
  • Locked the MZ, using the difference in the IR PDs as the error signal
  • Looked at the green difference for bounding "excess phase noise"
  • The spectra taken showed "glitchiness" and to get a low noise spectrum with ~5 averages, we needed to take ~10+ spectra. (Weird transient glitches)
  • Obtained the above linked plot for ~4 or 5 averages.

Calibration from this night

  • Green: [min,max] = [-4.7,7.2] = 11.9 Vpp (PDs Very Unbalanced)
  • IR: [min,max] = [-8,8.5]  = 16.5 Vpp (PDs Less Unbalanced)

After that, the thinking was to solder everything together and eliminate the general crappiness of the stuffable breadboard (crazy glitchy pickup). The frontend was still in some broken state

  • I took some old NIM box from our bridge subbasement E lab and gutted it.
  • I took about a week to make the circuit for the PD readout (slightly modifying crappy things as I went) - I will make this a sub entry later -
  • The circuit included
    • Green transimpedance amplifiers
    • PD Balancing via Potentiometers
    • Sum and Differencing circuits for both wavelengths
    • Mach Zehnder locking servo circuit
    • ISS servo circuit
    • Regulators
  • I also added surgical rubber hosing below the box to try to provide a better air seal
    • I am somewhat close to the aperture on my plastic box now - maybe a few beam widths
    • I also haven't done any blocking of where the wires go into the box, and may need to stuff this with some sort of foam
  • AT SOME POINT IN TAKING THE BOX ON AND OFF I MESSED UP THE ALIGNMENT INTO THE MACH ZEHNDER: I SAW THE KNOB ON AN INPUT MIRROR VISIBLY MOVE
  • I spent maybe 10 minutes playing with the knob and did some crude qualitative alignment into the ovens,
  • I need to do something much more quantitative b/c one oven is damaged (linkback): loss / beam quality / efficiency is rather sensitive to exactly what optical axis I'm using through the crystal

I then tried to put everything together and get a phase noise spectrum. Koji helped me figure out some set of things I was doing wrong.

  • I realigned the Mach Zehnder (and was able to set the absolute phase difference between the two wavelengths via translational movement of the beam axis)
  • The above is not obviously "bad" to do because there is already some trade off between contrast in each wavelength (due to angular dispersion), so it's just something to consider in my choice of (mis)alignment
  • I balanced the PDs by balancing the extrema which the Mach Zehnder difference signals swept through (was able balance to ~ 1%)
  • I locked the IR difference, and suppressed the error signal
  • I looked at the green difference, and it was not suppressed (Plot number 1)
  • The phase noise I saw in the green was inconsistent with the first plot linked (before I took everything apart to solder it together)

Calibration from this night I AM NOT POSITIVE THAT I DID NOT CHANGE THE GREEN TRANSIMPEDANCE RESISTOR (it is now 100k), and my balancing effects this up to 30%

  • Green: [min,max] = [-3.5,+3.5] = 7 Vpp (59% of previous calibration voltage)
  • IR: [min,max] = [-3.6,3.6]  = 7.2 Vpp (44% of previous calibration voltage - these TIAs are the same)

I Currently think that I increased the scatter in my ovens, and I am IR scatter limited on the green PDs. More to come.

I aso Adjusted the Power into the Mach Zehnder

 

Attachment 1: plottedApr20Phase.pdf  98 kB  Uploaded Wed May 19 20:26:09 2010  | Hide | Hide all
plottedApr20Phase.pdf
ELOG V3.1.3-